OpenStudy (simmy2497):

Is this the right answer to this question? (y-2=1(x-1) Write the point-slope form of the equation of the horizontal line that passes through the point (2, 1). Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution.

OpenStudy (tkhunny):

1) You have x and y backwards. 2) Horizontal lines don't have an x-component. 3) Horizontal lines have a slope of zero (0). You made up the one (1). Give it another go.

OpenStudy (simmy2497):

y=x-2?

OpenStudy (tkhunny):

Clarify #1 (2,1) means x = 2 and y = 1. That's what you have backwards. The slope of a horizontal line is ZERO (0). You keep using ONE (1). Use zero (0).

OpenStudy (simmy2497):

y-2=0(x-0) confused..

OpenStudy (tkhunny):

y-1 = 0*(x-2)

OpenStudy (tkhunny):

Use x and y properly. Use the correct slope. No need to be confused.

OpenStudy (simmy2497):

ohh ok thank you!

OpenStudy (tkhunny):

y = 1 No x-component.

OpenStudy (tkhunny):

{sigh} y = 1. There is no x in there. If you find an 'x' in there, it's not a horizontal line. Try not to confuse the OP. If we wish to talk about set definition, sure, we can do that. Again, try not to confuse the OP who only needs to understand that the slope is 0 and there will be no 'x' found in the formula.

OpenStudy (tkhunny):

We're talking about simple definitions, here, not theoretical constructs. Leave it alone. The actual characters used in the equation contain no "x". It needn't be any more difficult than that.